OpenStudy (anonymous):

Solid Mensuration: A regular hexagon A has the midpoints of its edges joined to form a smaller hexagon B. This process is repeated by joining the midpoints of the edges of hexagon B to get a third hexagon C. What is the ratio of the area of hexagon C to the area of hexagon A?

OpenStudy (compassionate):

\(\Huge\bf Welcome~to~ \color{#00B4ff}{Open}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge\bf\color{#7cc517}{Study~!!}~\) \[x^2=2(\frac{n}{2})^2-2(\frac{n}{2})^2~\cos(\alpha)\] \[Ratio = \frac{ 9 }{ 16}\] \[p = .75n\] \[k = .75^2\]

OpenStudy (anonymous):

thank you